# Weighted Fisher divergence for high-dimensional Gaussian variational inference

### v1

@@ Abstract
Bayesian inference has many advantages for complex models. However, standard Monte Carlo methods for summarizing the posterior can be computationally demanding, and it is attractive to consider optimization-based variational approximations. Our work considers Gaussian approximations with sparse precision matrices which are tractable to optimize in high-dimensional problems. Although the optimal Gaussian approximation is usually defined as the one closest to the target posterior in Kullback-Leibler divergence, it is useful to consider other divergences when the Gaussian assumption is crude, in order to capture important features of the posterior for a given application. Our work studies the weighted Fisher divergence, which focuses on gradient differences between the target posterior and its approximation, with the Fisher and score-based divergences being special cases. We make three main contributions. First, we compare approximations for weighted Fisher divergences under mean-field assumptions for both Gaussian and non-Gaussian targets with Kullback-Leibler approximations. Second, we go beyond mean-field and consider approximations with sparse precision matrices reflecting posterior conditional independence structure for hierarchical models. Using stochastic gradient descent to enforce sparsity, we develop two approaches to minimize the weighted Fisher divergence, based on the reparametrization trick and a batch approximation of the objective. Finally, we examine the performance of our methods for examples involving logistic regression, generalized linear mixed models and stochastic volatility models.
Fisher divergence,
Score-based divergence,
Stochastic gradient descent,
@@ 1 Introduction
Bayesian inference is a powerful tool for quantifying uncertainty, but it is demanding to implement for two reasons. Firstly, specifying a full probabilistic model for all unknowns and observables requires careful thought, and the components of the model need to be checked against the data. Secondly, Bayesian computations are difficult, requiring approximation of high-dimensional integrals. For many Bayesian models, exact posterior inference is infeasible, and a variety of numerical methods for summarizing the posterior are used in practice, such as Markov chain Monte Carlo (MCMC) and variational inference (VI). MCMC is asymptotically unbiased, which means that we can estimate posterior quantities as precisely as we wish with a large enough number of iterations. While MCMC is often treated as the gold standard for posterior estimation, its computational cost can be prohibitively high for large datasets or complex models (Robert and Casella, 2004; Maclaurin and Adams, 2015). On the other hand, VI reformulates posterior computation into an optimization problem by minimizing a divergence between the true posterior and a simpler variational distribution. This enables faster and more scalable inference, leveraging advances in optimization algorithms (Blei et al., 2017). As a result, VI is increasingly popular for its computational efficiency in large-scale problems.
@@
The performance of VI is largely determined by the family of variational approximations chosen, optimization technique, and divergence characterizing discrepancy between the true posterior and variational density. Much of the VI literature has focused on improving expressiveness of the variational family and enhancing optimization methods, often using Kullback-Leibler divergence (KLD) as the measure of approximation quality. To better capture the dependence structure among variables, which can be especially strong in hierarchical models, structured variational approximations that mimic the true dependency structure can be employed (Hoffman and Blei, 2015; Tan and Nott, 2018; Durante and Rigon, 2019; Tan, 2021). More recently, flow-based methods which transform an initial simple distribution into more flexible forms through a series of invertible transformations have been introduced (Rezende et al., 2014; Dinh et al., 2017; Agrawal and Domke, 2024). These approaches allow VI to capture highly complex posterior distributions, significantly enhancing the flexibility of the inference.
@@
Despite the popularity of KLD, studying alternatives can be important, particularly when using simple variational families which may be employed for tractability in high-dimensional problems. These approximations may not be capable of matching the target posterior closely, and choosing an appropriate divergence can help to capture the most important features of the posterior for a given application. A family of divergences including KLD as a special case is the Rényi’s $\alpha$ family (Li and Turner, 2016), with parameter $\alpha$ that can be adjusted to give Hellinger distance ($\alpha=0.5$), $\chi^{2}$-divergence ($\alpha=2$) and KLD ($\alpha=1$). The value of $\alpha$ can be chosen to balance between mode-seeking and mass-covering behavior, but when $\alpha\neq 1$, the most practical methods for optimizing the variational Rényi bound use biased stochastic gradients. Stein divergence has also emerged as a powerful objective for VI. Ranganath et al. (2016) introduced operator variational inference, a minimax approach that optimizes Stein discrepancies by constructing variational objectives based on Stein operators. Liu and Wang (2016) developed Stein variational gradient descent, which uses kernelized Stein discrepancies to iteratively transform particles toward the posterior. In this article, we explore use of the weighted Fisher divergence in Gaussian VI, focusing on the Fisher and score-based divergences as special cases. The definitions and motivations for studying these divergences are presented below.
@@ 1.1 Weighted Fisher divergence
Let ${\text{E}}_{q}$ denote expectation with respect to $q(\theta)$. As $\log p(y)=\text{\rm KL}(q\|p)+\mathcal{L}$, where $\mathcal{L}={\text{E}}_{q}\{\log p(y,\theta)-\log q(\theta)\}$, minimizing the KLD is equivalent to maximizing an evidence lower bound $\mathcal{L}$ on $\log p(y)$, which does not depend on the normalizing constant of the true posterior.
@@
[…]
$={\text{E}}_{q}\{\|\nabla_{\theta}\log q(\theta)-\nabla_{\theta}\log p(\theta|y)\|_{M}^{2}\},$
@@
where $\|\cdot\|_{M}$ is the $M$-weighted vector norm defined as $\|z\|_{M}=\sqrt{z^{\top}Mz}$ and $M$ is a positive semi-definite matrix. Like KLD, $S_{M}(q\|p)$ is asymmetric, non-negative, and vanishes when $q(\theta)=p(\theta|y)$. Let $h(\theta)=p(y|\theta)p(\theta)$. Then $\nabla_{\theta}\log p(\theta|y)=\nabla_{\theta}\log h(\theta)$, which is independent of the unknown normalizing constant $p(y)$. Similarly, if the variational density contains an unknown normalizing constant, this does not need to be computed to evaluate the weighted Fisher divergence. Unlike the evidence lower bound, the weighted Fisher divergence provides a direct measure of the distance between the true posterior and variational density.
@@
When $M$ is the identity matrix $I$, $S_{I}(q\|p)$ is known as the Fisher divergence (FD, Hyvärinen, 2005), denoted hereafter as $F(q\|p)$. When $q(\theta)$ is ${\text{N}}(\mu,\Sigma)$ and $M$ is its covariance matrix $\Sigma$, $S_{\Sigma}(q\|p)$ is referred to as score-based divergence (SD) in Cai et al. (2024), denoted as $S(q\|p)$ henceforth. Cai et al. (2024) derived closed-form updates for Gaussian variational parameters in a batch and match (BaM) algorithm based on score-based divergence, and showed that $S(q\|p)$ is affine invariant while $F(q\|p)$ is not. This means that $S(\tilde{q}\|\tilde{p})=S(q\|p)$ if $\tilde{p}$ and $\tilde{q}$ denote the densities of $p$ and $q$ respectively after applying an affine transformation of $\theta$.
@@
In recent years, there is increasing interest in use of the weighted Fisher divergence in VI. Huggins et al. (2020) showed that the Fisher divergence defined in terms of the generalized $\ell_{p}$ norm is an upper bound to the $p$-Wasserstein distance, and its optimization ensures closeness of the variational density to the true posterior in terms of important point estimates and uncertainties. Yang et al. (2019) derived an iteratively reweighted least squares algorithm for minimizing the Fisher divergence in exponential family based variational approximations. Elkhalil et al. (2021) employed the factorizable polynomial exponential family as variational approximation in their Fisher autoencoder framework, which has performance competitive with existing ones. Modi et al. (2023) developed Gaussian score matching variational inference with closed form updates, by minimizing the KLD between a target and Gaussian variational density subject to a matching score function constraint. For implicit variational families structured hierarchically, Yu and Zhang (2023) used the Fisher divergence to reformulate the optimization objective into a minimax problem. Cai et al. (2024) proposed a variational family built on orthogonal function expansions, and transformed the optimization objective into a minimum eigenvalue problem using Fisher divergence.
@@
Our contributions in this article are threefold. First, we study behavior of the weighted Fisher divergence in mean-field Gaussian VI for both Gaussian and non-Gaussian targets, showing its tendency to underestimate the posterior variance more severely than KLD. Second, we develop Gaussian VI for high-dimensional hierarchical models where posterior conditional independence structure is captured via a sparse precision matrix. Sparsity is enforced by computing updates using stochastic gradient descent (SGD) and two approaches are proposed for minimizing the weighted Fisher divergence. The first relies on the reparametrization trick (Kingma and Welling, 2014), while the second relies on a biased estimate of the objective based on a batch of samples at each iteration (Elkhalil et al., 2021; Cai et al., 2024). Third, we study the variance of unbiased gradient estimates computed using the reparametrization trick and limiting behavior of the batch approximated weighted Fisher divergence under the mean-field assumption.
@@
This article is organized as follows. We study the quality of posterior mode and variance approximations for Gaussian and non-Gaussian targets in Sections 2 and 3 respectively, when using the weighted Fisher divergence in VI. Section 4 introduces Gaussian VI for hierarchical models by using a sparse precision matrix to capture posterior conditional independence. Two SGD approaches for minimizing the weighted Fisher divergence are proposed in Sections 5 and 6, based respectively on the reparametrization trick and batch approximation. Experimental results are discussed in Section 7 with applications to logistic regression, generalized linear mixed models (GLMMs) and stochastic volatility models. Section 8 concludes the paper with a discussion of the findings and future work.
2 Ordering of divergences for
Gaussian target
@@ 2 Ordering of divergences for
Accurate estimation of the posterior variance is important in VI, as it affects uncertainty quantification in Bayesian inference. Here, we establish an ordering of the weighted Fisher and KL divergences according to the estimated posterior variance when the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$ with a non-diagonal precision matrix $\Lambda$. All divergences considered can recover the true mean $\nu$ and precision $\Lambda$ when the variational family is also Gaussian with a full covariance matrix. However, the computational cost of optimizing a full-rank Gaussian variational approximation can be prohibitive for high-dimensional models. A widely used alternative is the mean-field Gaussian variational approximation, $q(\theta)={\text{N}}(\mu,\Sigma)$, with a diagonal covariance matrix $\Sigma$. The mean-field assumption simplifies the optimization but tends to underestimate the true posterior variance under the KLD (Blei et al., 2017; Tan and Nott, 2018; Giordano et al., 2018). Here, we examine the severity of posterior variance underestimation under the weighted Fisher divergence compared to KLD.
@@
Our results in this section generalize similar results in Margossian et al. (2024) from score-based divergences to the more general class of weighted Fisher divergences. For KLD under the mean-field assumption, Margossian et al. (2024) showed that the posterior mean can be recovered ($\mu=\nu$) and the optimal variance parameter is
$\Sigma^{\text{KL}}_{ii}=\Lambda_{ii}^{-1}\quad\text{for}\quad i=1,\dots,d.$
@@
Thus, the precision is matched by the variational density, but the variance is underestimated. Lemma 1 presents the weighted Fisher divergence for a general weight matrix $M$, which is $I_{d}$ in the Fisher divergence and $\Sigma$ in the score-based divergence.
@@
Proof.
The $M$-weighted Fisher divergence is
$=\text{\rm tr}\{(\Sigma^{-1}-\Lambda)M(\Sigma^{-1}-\Lambda)\Sigma\}$
$\quad+(\mu-\nu)^{\top}\Lambda M\Lambda(\mu-\nu)$
$\quad+(\mu-\nu)^{\top}\Lambda M\Lambda(\mu-\nu).$
The final result arises from $\text{\rm tr}(AB)=\sum_{i=1}^{d}A_{ii}B_{ii}$ if $A$ is a diagonal matrix. ∎
From Lemma 1, $\nabla_{\mu}S_{M}(q\|p)=2\Lambda M\Lambda(\mu-\nu)$. Thus, $\nabla_{\mu}S_{M}(q\|p)=0$ implies $\mu=\nu$, and the true posterior mean is recovered for any $M$-weighted Fisher divergence where $M$ is independent of $\mu$. Under the mean-field assumption, at this optimal value of $\mu$,
@@
$\Sigma_{ii}=\sqrt{M_{ii}/(\Lambda M\Lambda)_{ii}}\quad\text{for}\quad i=1,\dots,d.$
@@
$\Sigma_{ii}=\sqrt{\frac{M_{ii}}{\sum_{j=1}^{d}M_{jj}\Lambda_{ij}^{2}}}\quad\text{for}\quad i=1,\dots,d.$ (1)
When $M_{ii}=1$ $\forall$ $i$, we recover the Fisher divergence for which the optimal variance parameters are
$\Sigma^{\text{F}}_{ii}=\frac{1}{\sqrt{\sum_{j=1}^{d}\Lambda_{ij}^{2}}}\quad\text{for}\quad i=1,\dots,d.$ (2)
Optimal variational parameters for score-based divergence under mean-field assumption have been presented in Margossian et al. (2024), and a discussion is included here for completeness. Plugging $M=\Sigma$ in Lemma 1,
@@
at the optimal value of $\mu$. Let $\mathbf{s}=(s_{1},\dots,s_{d})^{\top}$ such that $s_{i}=\Sigma_{ii}\Lambda_{ii}\geq 0$, and $H$ be a $d\times d$ symmetric matrix with $H_{ij}=\Lambda_{ij}^{2}/(\Lambda_{ii}\Lambda_{jj})$. Then $S(q\|p)=d+2F(\mathbf{s})$, where
@@
Thus the optimal $\Sigma^{\text{S}}_{ii}$ that minimizes $S(q\|p)$ can be obtained by solving a non-negative quadratic program (NQP) for $\mathbf{s}$. NQP is the problem of minimizing the quadratic objective function in (3) subject to the constraint $s_{i}\geq 0$ $\forall$ $i$. Since $\Lambda$ is positive definite,
@@
for any $x=(x_{1},\dots,x_{d})^{\top}\in\mathbb{R}^{d}$ and $y=(x_{1}/\Lambda_{11},\dots,x_{d}/\Lambda_{dd})^{\top}$. Thus $H$ is symmetric positive definite, which implies that $F(\mathbf{s})$ is bounded below and its optimization is convex. However, there is no analytic solution for the global minimum due to the non-negativity constraints and iterative solutions are required (Sha et al., 2003). The Karush-Kuhn-Tucker (KKT) conditions are first derivative tests that can be used to check whether a solution returned by an iterative solver is indeed a local optimum. For the NQP in (3), the KKT conditions state that $\forall$ $i=1,\dots,d$, either (a) $s_{i}=0$ and $(Hs)_{i}>1$ or (b) $s_{i}>0$ and $(Hs)_{i}=1$. Note that $\nabla_{\mathbf{s}}F(\mathbf{s})=H\mathbf{s}-{\bf 1}$. These conditions correspond to cases where the constraint is active or inactive at the optimum. Case (a) implies $\Sigma_{ii}^{\text{S}}=0$, meaning that the variational density collapses to a point estimate in the $i$th dimension. Note that the KL and Fisher divergences do not face this issue of “variational collapse”. Case (b) implies
@@
$\implies\sum_{j=1}^{d}\Lambda_{ij}^{2}\Sigma_{jj}=\Lambda_{ii}.$
Figure 1: Variance parameter comparisons for Fisher and score-based divergence.
@@
Suppose the target is a multivariate Gaussian with non-diagonal precision matrix $\Lambda$, and the variational family is Gaussian with diagonal covariance matrix $\Sigma$. Let $\Sigma^{\text{KL}}_{ii}$, $\Sigma^{\text{M}}_{ii}$ and $\Sigma^{\text{S}}_{ii}$ denote the optimal value of the $i$th diagonal element of $\Sigma$ obtained by minimizing the KL, $M$-weighted Fisher and score-based divergences respectively, where $M$ is a positive definite diagonal matrix independent of $\Sigma$. Then
$\Sigma^{\text{M}}_{ii}\leq\Sigma^{\text{KL}}_{ii}\quad\text{and}\quad\Sigma^{\text{S}}_{ii}\leq\Sigma^{\text{KL}}_{ii}\quad\text{for}\quad i=1,\dots,d.$
Moreover, $\exists\,i\in\{1,\dots,d\}$ such that $\Sigma_{ii}^{M}<\Sigma_{ii}^{KL}$ and $\Sigma_{ii}^{S}<\Sigma_{ii}^{KL}$.

### v2

@@ Abstract
Bayesian inference has many advantages for complex models, but standard Monte Carlo methods for summarizing the posterior can be computationally demanding, and it is attractive to consider optimization-based variational methods. Our work considers Gaussian approximations with sparse precision matrices which are tractable to optimize in high-dimensions. The optimal Gaussian approximation is usually defined as being closest to the posterior in Kullback-Leibler divergence, but it is useful to consider other divergences when the Gaussian assumption is crude, to capture important posterior features for given applications. Our work studies the weighted Fisher divergence, which focuses on gradient differences between the target posterior and its approximation, with the Fisher and score-based divergences as special cases. We make three main contributions. First, we compare approximations for weighted Fisher divergences under mean-field assumptions for Gaussian and non-Gaussian targets with Kullback-Leibler approximations. Second, we go beyond mean-field and consider approximations with sparse precision matrices reflecting posterior conditional independence structure for hierarchical models. Using stochastic gradient descent to enforce sparsity, we develop two approaches to minimize the Fisher and score-based divergences, based on the reparametrization trick and a batch approximation of the objective. Finally, we study the performances of our methods using logistic regression, generalized linear mixed models and stochastic volatility models.
@@ 1 Introduction
Bayesian inference is a powerful tool for quantifying uncertainty, but it is demanding to implement for two reasons. First, specifying a full probabilistic model for all unknowns and observables requires careful thought, and components of the model need to be checked against the data. Second, Bayesian computations are difficult, requiring approximation of high-dimensional integrals. For many Bayesian models, exact posterior inference is infeasible, and a variety of numerical methods for summarizing the posterior are used in practice, such as Markov chain Monte Carlo (MCMC) and variational inference (VI). MCMC is often asymptotically unbiased, in that we can estimate posterior quantities as precisely as we wish with a large enough number of iterations, although certain variants (e.g. non-reversible methods) may incur a small bias. While MCMC is often treated as the gold standard for posterior estimation, its computational cost can be prohibitively high for large datasets or complex models (Robert and Casella, 2004; Maclaurin and Adams, 2015). On the other hand, VI reformulates posterior approximation into an optimization problem by minimizing a divergence between the true posterior and a simpler variational distribution. This enables faster and more scalable inference, leveraging advances in optimization algorithms (Blei et al., 2017). As a result, VI is increasingly popular for its computational efficiency in large-scale problems.
@@
The performance of VI is largely determined by the family of variational approximations chosen, optimization technique, and divergence characterizing discrepancy between the true posterior and variational density. Much of the VI literature has focused on improving expressiveness of the variational family and enhancing optimization methods, often using Kullback-Leibler divergence (KLD) as a measure of approximation quality. To better capture dependence structure among variables, which can be especially strong in hierarchical models, partially factorized VI (Goplerud et al., 2025) or structured variational approximations that mimic the true dependency structure (Hoffman and Blei, 2015; Tan and Nott, 2018; Durante and Rigon, 2019; Tan, 2021) can be employed. More recently, flow-based methods which transform an initial simple distribution into more flexible forms through a series of invertible transformations have been introduced (Rezende et al., 2014; Dinh et al., 2017; Agrawal and Domke, 2024). These approaches allow VI to capture highly complex posterior distributions, significantly enhancing the flexibility of the inference.
@@
Despite the popularity of KLD, studying alternatives is important, particularly when using simple variational families which may be employed for tractability in high-dimensional problems. These approximations may not be capable of matching the posterior closely, and choosing an appropriate divergence can help to capture the most important features of the posterior for a given application. A family of divergences including KLD as a special case is the Rényi’s $\alpha$ family (Li and Turner, 2016), where $\alpha$ can be adjusted to give Hellinger distance ($\alpha=0.5$), $\chi^{2}$-divergence ($\alpha=2$) and KLD ($\alpha=1$). While $\alpha$ can help to balance between mode-seeking and mass-covering behavior, the most practical methods for optimizing the variational Rényi bound use biased stochastic gradients when $\alpha\neq 1$. Stein divergence has also emerged as a powerful objective for VI. Ranganath et al. (2016) introduced operator variational inference, a minimax approach that optimizes Stein discrepancies by constructing variational objectives based on Stein operators. Liu and Wang (2016) developed Stein variational gradient descent, which uses kernelized Stein discrepancies to iteratively transform particles toward the posterior. In this article, we explore use of the weighted Fisher divergence in Gaussian VI, focusing on the Fisher and score-based divergences as special cases. The definitions and motivations for studying these divergences are presented below.
@@ 1.1 Weighted Fisher divergence
Let ${\text{E}}_{q}$ denote expectation with respect to $q(\theta)$. As $\log p(y)=\text{\rm KL}(q\|p)+\mathcal{L}$, where $\mathcal{L}={\text{E}}_{q}\{\log p(y,\theta)-\log q(\theta)\}$, minimizing the KLD is equivalent to maximizing an evidence lower bound $\mathcal{L}$ on $\log p(y)$, which does not depend on normalizing constant of the true posterior.
@@
$S_{M}(q\|p)$ $=\int q(\theta)\bigg\|\nabla_{\theta}\log\frac{q(\theta)}{p(\theta|y)}\bigg\|_{M}^{2}d\theta,$
@@
where $\|\cdot\|_{M}$ is the $M$-weighted vector norm defined as $\|z\|_{M}=\sqrt{z^{\top}Mz}$ and $M$ is a positive semi-definite matrix. Like KLD, $S_{M}(q\|p)$ is asymmetric, non-negative, and vanishes when $q(\theta)=p(\theta|y)$. Let $h(\theta)=p(y|\theta)p(\theta)$. Then $\nabla_{\theta}\log p(\theta|y)=\nabla_{\theta}\log h(\theta)$, which is independent of the unknown normalizing constant $p(y)$. Similarly, if $q(\theta)$ contains an unknown normalizing constant, this is not required to evaluate the weighted Fisher divergence. Unlike the evidence lower bound, the weighted Fisher divergence provides a direct measure of the distance between the true posterior and variational density.
@@
When $M$ is the identity matrix $I$, $S_{I}(q\|p)$ is known as Fisher divergence (FD, Hyvärinen, 2005), denoted hereafter as $F(q\|p)$. When $q(\theta)$ is ${\text{N}}(\mu,\Sigma)$ and $M$ is its covariance matrix $\Sigma$, $S_{\Sigma}(q\|p)$ is known as score-based divergence (SD) in Cai et al. (2024), denoted as $S(q\|p)$ henceforth. Cai et al. (2024) derived closed-form updates for Gaussian variational parameters in a batch and match (BaM) algorithm based on the SD, and showed that $S(q\|p)$ is affine invariant while $F(q\|p)$ is not. This means that $S(\tilde{q}\|\tilde{p})=S(q\|p)$ if $\tilde{p}$ and $\tilde{q}$ denote the densities of $p$ and $q$ respectively after an affine transformation of $\theta$.
In sliced score matching (Song et al., 2020), the scores are projected onto randomly generated vectors $v$ before comparison for dimension reduction, and the weight matrix $M={\text{E}}(vv^{\top})$. Liu et al. (2022) applied the weighted Fisher divergence in estimating the parameters of truncated densities, whose normalizing constants are intractable, and the weight function is the shortest distance between a data point and the boundary of the domain. The weighted Fisher divergence is also widely used in training score-based generative models (Song et al., 2021), where a forward diffusion and reverse-time process are defined through stochastic differential equations (SDEs). The scores are estimated via neural networks and trained using a time integrated weighted Fisher divergence, where the weight matrix depends on a function of time specified in the SDE (Huang et al., 2021; Lu et al., 2022). The above choices of $M$ are not directly applicable or lack intrinsic motivation in our setting, and hence we focus primarily on the FD and SD, as they represent natural and widely studied choices in VI. However, our results in Section 2 also consider general constant weight matrices $M$, besides the FD and SD.
@@
In recent years, there is increasing interest in use of the weighted Fisher divergence in VI. Huggins et al. (2020) showed that the Fisher divergence defined in terms of the generalized $\ell_{p}$ norm is an upper bound to the $p$-Wasserstein distance, and its optimization ensures closeness of the variational density to the true posterior in terms of important point estimates and uncertainties. Yang et al. (2019) derived an iteratively reweighted least squares algorithm for minimizing the FD in exponential family based variational approximations, while Elkhalil et al. (2021) employed the factorizable polynomial exponential family as variational approximation in their Fisher autoencoder framework. Modi et al. (2023) developed Gaussian score matching variational inference with closed form updates, by minimizing the KLD between a target and Gaussian variational density subject to a matching score function constraint. For implicit variational families structured hierarchically, Yu and Zhang (2023) used the FD to reformulate the optimization objective into a minimax problem. Cai et al. (2024) proposed a variational family built on orthogonal function expansions, and transformed the optimization objective into a minimum eigenvalue problem using the FD.
@@
Our contributions in this article are fourfold. First, we study behavior of the weighted Fisher divergence in mean-field Gaussian VI for Gaussian and non-Gaussian targets, showing its tendency to underestimate the posterior variance more severely than KLD. Second, we develop Gaussian VI for high-dimensional hierarchical models for which posterior conditional independence structure is captured via a sparse precision matrix. Sparsity is enforced by using stochastic gradient descent (SGD), and two distinct approaches are proposed for minimizing the FD and SD. Algorithms based on unbiased gradients computed using the reparameterization trick (Kingma and Welling, 2014) are denoted as FDr and SDr (“r” for reparameterization trick), while algorithms that rely on a batch approximation of the objective at each iteration are denoted by FDb and SDb (“b” for batch approximation). Third, we study the variance of unbiased gradient estimates computed using the reparametrization trick, and limiting behavior of the batch approximated FD and SD under mean-field. Finally, we present extensive experiments demonstrating that methods based on the reparameterization trick (FDr and SDr) suffer from high variations in gradients and perform poorly relative to baselines such as KLD and BaM. In contrast, methods based on the batch approximation (FDb and SDb) converge more rapidly and scale more efficiently to high-dimensional models.
@@
This article is organized as follows. We study the quality of posterior mean, mode and variance approximations for Gaussian and non-Gaussian targets in Sections 2 and 3 respectively, when using the weighted Fisher divergence in VI. Section 4 introduces Gaussian VI for hierarchical models by capturing posterior conditional independence via a sparse precision matrix. Two SGD approaches for minimizing the weighted Fisher divergence are proposed in Sections 5 and 6, based respectively on the reparametrization trick and batch approximation. Experimental results are discussed in Section 7 with applications to logistic regression, generalized linear mixed models (GLMMs) and stochastic volatility models. Section 8 concludes the paper with a discussion.
2 Ordering of divergences for Gaussian target
@@
Accurate estimation of the posterior variance is important in VI, as it affects uncertainty quantification in Bayesian inference. Here, we establish an ordering of the weighted Fisher and KL divergences according to the estimated posterior variance when the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$ with a precision matrix $\Lambda$. All divergences considered can recover the true mean $\nu$ and precision matrix $\Lambda$ when the variational family is also Gaussian with a full covariance matrix. However, the computation cost of optimizing a full-rank Gaussian variational approximation can be prohibitive for high-dimensional models. A widely used alternative is the mean-field Gaussian variational approximation, $q(\theta)={\text{N}}(\mu,\Sigma)$, with a diagonal covariance matrix $\Sigma$. The mean-field assumption simplifies the optimization but tends to underestimate the true posterior variance under KLD (Blei et al., 2017; Tan and Nott, 2018; Giordano et al., 2018). Here, we examine the severity of posterior variance underestimation under the weighted Fisher divergence compared to KLD.
@@
Our results in this section generalize similar results in Margossian et al. (2024) from SD to the general class of weighted Fisher divergences. For KLD under the mean-field assumption, Margossian et al. (2024) showed that the posterior mean can be recovered ($\hat{\mu}=\nu$) and the optimal variance parameter is
$\hat{\Sigma}^{\text{KL}}_{ii}=1/\Lambda_{ii}\quad\text{for}\quad i=1,\dots,d.$
@@
Thus, the precision is matched by the variational density, but the variance is underestimated. Lemma 1 presents the weighted Fisher divergence for a general weight matrix $M$, which is $I_{d}$ in FD and $\Sigma$ in SD.
@@
From Lemma 1, $\nabla_{\mu}S_{M}(q\|p)=2\Lambda M\Lambda(\mu-\nu)$. Thus, $\nabla_{\mu}S_{M}(q\|p)=0$ implies $\hat{\mu}=\nu$, and the true posterior mean is recovered for any $M$-weighted Fisher divergence where $M$ is independent of $\mu$. Under the mean-field assumption, at this optimal value $\hat{\mu}$,
@@
$\hat{\Sigma}_{ii}=\sqrt{M_{ii}/(\Lambda M\Lambda)_{ii}}\quad\text{for}\quad i=1,\dots,d.$
@@
$\hat{\Sigma}_{ii}=\sqrt{\frac{M_{ii}}{\sum_{j=1}^{d}M_{jj}\Lambda_{ij}^{2}}}\quad\text{for}\quad i=1,\dots,d.$ (1)
When $M_{ii}=1$ $\forall$ ii, we recover the FD for which the optimal variance parameters are
$\hat{\Sigma}^{\text{F}}_{ii}=\frac{1}{\sqrt{\sum_{j=1}^{d}\Lambda_{ij}^{2}}}\quad\text{for}\quad i=1,\dots,d.$ (2)
Optimal variational parameters for SD under the mean-field assumption have been presented in Margossian et al. (2024), and a discussion is included here for completeness. Plugging $M=\Sigma$ in Lemma 1,
@@
at the optimal value $\hat{\mu}$. Let $\mathbf{s}=(s_{1},\dots,s_{d})^{\top}$ such that $s_{i}=\Sigma_{ii}\Lambda_{ii}\geq 0$, and $H$ be a $d\times d$ symmetric matrix with $H_{ij}=\Lambda_{ij}^{2}/(\Lambda_{ii}\Lambda_{jj})$. Then $S(q\|p)=d+2F(\mathbf{s})$, where
@@
Thus the optimal $\hat{\Sigma}^{\text{S}}_{ii}$ that minimizes $S(q\|p)$ can be obtained by solving a non-negative quadratic program (NQP) for $\mathbf{s}$. NQP is the problem of minimizing the quadratic objective function in (3) subject to the constraint $s_{i}\geq 0$ $\forall$ ii. Since $\Lambda$ is positive definite,
@@
for any $x=(x_{1},\dots,x_{d})^{\top}\in\mathbb{R}^{d}$ and $y=(x_{1}/\Lambda_{11},\dots,x_{d}/\Lambda_{dd})^{\top}$. Thus $H$ is symmetric positive definite, which implies that $F(\mathbf{s})$ is bounded below and its optimization is convex. However, there is no analytic solution for the global minimum due to the non-negativity constraints and iterative solutions are required (Sha et al., 2003). The Karush-Kuhn-Tucker (KKT) conditions are first derivative tests that can be used to check whether a solution returned by an iterative solver is indeed a local optimum. For the NQP in (3), the KKT conditions state that $\forall$ $i=1,\dots,d$, either (a) $s_{i}=0$ and $(Hs)_{i}>1$ or (b) $s_{i}>0$ and $(Hs)_{i}=1$. Note that $\nabla_{\mathbf{s}}F(\mathbf{s})=H\mathbf{s}-{\bf 1}$. These conditions correspond to cases where the constraint is active or inactive at the optimum. Case (a) implies $\hat{\Sigma}_{ii}^{\text{S}}=0$, meaning that the variational density collapses to a point estimate in the $i$th dimension. Note that KLD and FD do not face this issue of “variational collapse”. Case (b) implies
@@
$\implies\sum_{j=1}^{d}\Lambda_{ij}^{2}\hat{\Sigma}_{jj}^{\text{S}}=\Lambda_{ii}.$
Figure 1: Variance parameter comparisons for FD and SD. The red, blue and green regions indicate where $\Sigma_{ii}^{S}\leq\Sigma_{ii}^{F}$ for all cases, only two cases and only one case respectively. The orange-bordered region indicates where $\Lambda$ is diagonally dominant.
@@
Suppose the target is a multivariate Gaussian with precision matrix $\Lambda$, and the variational family is Gaussian with diagonal covariance matrix $\Sigma$. Let $\hat{\Sigma}^{\text{KL}}_{ii}$, $\hat{\Sigma}^{\text{M}}_{ii}$ and $\hat{\Sigma}^{\text{S}}_{ii}$ denote the optimal value of the $i$th diagonal element of $\Sigma$ obtained by minimizing the KL, $M$-weighted Fisher and score-based divergences respectively, where $M$ is a positive definite diagonal matrix independent of $\Sigma$. Then
$\hat{\Sigma}^{\text{M}}_{ii}\leq\hat{\Sigma}^{\text{KL}}_{ii}\quad\text{and}\quad\hat{\Sigma}^{\text{S}}_{ii}\leq\hat{\Sigma}^{\text{KL}}_{ii}\quad\text{for}\quad i=1,\dots,d,$
and $\exists\,i\in\{1,\dots,d\}$ such that $\hat{\Sigma}_{ii}^{M}<\hat{\Sigma}_{ii}^{\text{KL}}$ and $\hat{\Sigma}_{ii}^{S}<\hat{\Sigma}_{ii}^{\text{KL}}$.